Calculate the number of milliliters of 0.618 M KOH required to precipitate all of the Cr3+ ions in 194 mL of 0.621 M CrCl3 solution as Cr(OH)3. The equation for the reaction is:


CrCl3(aq) + 3KOH(aq) Cr(OH)3(s) + 3KCl(aq)


1
Expert's answer
2021-10-07T03:45:03-0400

CrCl3(aq) + 3KOH(aq) "\to"  Cr(OH)3(s) + 3KCl(aq)

Mole ratio of CrCl3 and KOH is 1:3

Moles of Cr3+ = Molarity × Volume

= 0.621× 194 = 120.474 moles

Moles of KOH = 3× 120.474 = 361.422 moles

Hence Volume of KOH = moles of KOH/Molarity of KOH

= 361.422 moles/0.618M = 584.83 milliliters